SAAO Telescope Time Application^{SAAO, PO Box 9, Observatory 7935, South Africa}^{Contact: observer@saao.ac.za}  ^{Email: observer@saao.ac.za}  ^{SAAO Reference:}^{Date Stamp:}^{Ph: +27 21 447 0025}^{1.} ^{Names and Institutions of Applicants}  ^{Expected Observers}^{Principal Contact:}^{Telephone:}  ^{Postal Address:}^{Fax:}^{Email:}^{Are the observations required for the completion of a student’s thesis?}^{Name of Student(s):}  ^{Name of Supervisor(s):}^{2.} ^{Quarter}^{3.} ^{Title of Project}^{Summary of Scientific Objectives}^{4.} ^{Telescope, Instrument and Detector}^{Telescope:}  ^{Instrument(s):}  ^{Detector(s):}  ^{Gratings/Filters:}^{5.} ^{Scheduling Information}^{Number of weeks requested}^{this quarter}  ^{Dark:}  ^{Grey:}  ^{Bright:}^{Minimum useful allocation}  ^{Dark:}  ^{Grey:}  ^{Bright:}^{Additional weeks required to}^{complete project in future}  ^{Dark:}  ^{Grey:}  ^{Bright:}  ^{If grey, is 1st or}^{3rd qrt preferred}^{Preferred dates:}^{Impossible dates:}^{(Give reasons)}^{Special scheduling constraints (eg}^{clashes with other applications,}^{lunar position or quarter, daytime}^{observing, requirements if more than}^{one instrument has been requested)}

6. Observing Requirements Average signal-to-noise ratio required: Any seeing limitations: Required spectral resolution (if applicable): Range in Brightness/ Surface Brightness: 7. Assistance Required Night assistants are usually only available for the first few nights of runs on the 1.9-m telescope. Supervisors have to accompany inexperienced students. 8. A short description of your project for the public that may be used for public relations purposes during your observing period. 9. List of Principal Targets (a representative list) 10. Other applications for observing time for this project in this quarter Telescope/satellite:  Title of programme: (include applications from all telescopes) 11. Report on previous applications for time for this and related projects. Show all allocations of SAAO time in past 2 years SAAO Reference Allocation % useful Publications (or Comments) (give # if known) 12. Scientific Justification: Attach a case of no more than 2 pages, in a font no smaller than 11pt.

Name α δ mag Comment